[Hallc_running] [New Logentry] Hall Status and plan

ciprian at jlab.org ciprian at jlab.org
Tue Oct 12 08:40:03 EDT 2021

Logentry Text:
RC Daily Meetings are at 10 am, remote: https://bluejeans.com/861439920

Accelerator Status
# Setup for Hall A will started around 4PM yesterday and after some difficulties we received beam back around 9:30 last night. 
#* The setup will continue this swing shift when Hall A will lock up, although we should expect beam for most of their setup. We will probably have higher chances of trips due to FSDs originating in the Hall A line.

Hall C Status
# Stephen and Dave moved the HMS yesterday to 58.535 degreees (https://logbooks.jlab.org/entry/3922017)
#* We reduced our current to 42 uA for this configuration based on the rates we were seeing in the SHMS (https://logbooks.jlab.org/entry/3922033)
#* Ion chambers did not need to be recalibrated (https://logbooks.jlab.org/entry/3922036)
#* Amy locked out the SHMS rotation (https://logbooks.jlab.org/entry/3922027)
# So far we have collected approximately 280 pions at this new kinematic (https://logbooks.jlab.org/entry/3922645)

Current Run Plan
# Continue data taking until early Nov


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3922686
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20211012/99f5a690/attachment.html>

More information about the Hallc_running mailing list